Serena is nine years old. Her mother is 39 years old. In how many years will her mother be three times as old as Serena?
Answer by stanbon(75887)

Serena is nine years old. Her mother is 39 years old. In how many years will her mother be three times as old as Serena?
----
s = 9
m = 39
-------
Let the "number of years" asked for be "x".
-----
Equation:
39 + x = 3(9 + x)
-----
39 + x = 27 + 3x
2x = 12
x = 6 years
===============
Cheers,
Stan H.